Electronics engineering: FILTERS


Design a 2nd order crossover network with the following specifications:  

  • High Pass cutoff frequency = 3100 Hz
  • Low Pass cutoff frequency = 310 Hz
  • Bandpass frequency = 310Hz to 3100Hz   
1
Expert's answer
2021-05-31T06:39:54-0400

An Op-Amp based bandpass filter is as shown below




The transfer function "\frac{V_0(s)}{V_{in}(s)}=-\frac{Z_0(s)}{Z_{in}(s)}"

"\frac{V_0(s)}{V_{in}(s)}=-[\frac{R_2}{(1+sC_2R_2)}]*[\frac{sc_1}{(1+sC_1R_1)}]=\frac{sc_1R_2}{(1+s \tau_1)(1+s \tau_2)} \implies \tau =RC"

The circuit will behave as BPF if "\tau _1> \tau_2"

"f_L=\frac{1}{2 \pi R_1C_1}"

"f_H=\frac{1}{2 \pi R_2C_2}"

"R_1=\frac{1}{2 \pi *310*47*10^{-9}}=10.924k \Omega"

Commercial available resistance nearest to calculated R1 is 11k"\Omega"

"\implies f_L=\frac{1}{2 \pi R_1C_1} =\frac{1}{2 \pi 11*10^3*47*10^{-9}} =307.84 Hz"

Therefore, "R_1=11k \Omega; C_1=47nF"

"C_2=\frac{1}{2 \pi R_2f_H} =\frac{1}{2 \pi *3100*20*10^3} = 2.567 nF"

Commercial available capacitor nearest to calculated C2 is 2.7nF

"f_H=\frac{1}{2 \pi R_2C_2}=\frac{1}{2 \pi*20*10^3*2.7*10^{-9}}=2947.31 Hz"

Therefore, "R_2=20k \Omega; C_2=2.7nF"